A) You are planning an experiment using the monolayer cell line X. You know that 2.70E+05 X cells gives you a 90.00% confluency in a T25 flask. Determine the number of cells needed to add to a T75 flask to obtain 60.00% confluency. Please answer using the scientific notation and one decimal point.

T25 flask area 25cm^2, volume 5ml

T75 flask area 75cm^2, volume 20ml

B) . If X cells are resuspended in media at a concentration of 7000 cells/mL, how much of the cell suspension will you need to add to the new flask to get the number of cells you calculated in part a.? Please answer using two significant digits.

edit: same cell count used in T25
Answer #1

Confluency refers to the percentage of area of the culture disc covered by adherent cells.

T25 flask area = 25 cm^2

No. of cells = 2.70E+05 cells

Culture is 90% confluent, this means 90% of 25 cm^2 is covered by cells

Area covered by cells = 90/100 *25 = 22.5 cm^2 = 22.5E+08 µm^2

Since it is a monolayer culture, 2.70E+05 cells are covering area of 22.5E+08 µm^2

Area of one cell = (22.5 x 10^8)/(2.7 x 10^5) = 8.33 x 10^3 µm^2

                              = 8330 µm^2

For T75 flask, area = 75 cm^2, confluency required = 60%

Area to be covered by cells for 60% confluency = (60/100) x 75 = 45 cm^2 = 45 x 10^8 µm^2

No. of the cells that would cover this area = (Area needed for 60% confluency)/Area of one cell

               = 45 x 10^8/8330 = 0.0054 x 10^8 = 5.4 x 10^5 cells

Cells to be added in T75 flask = (5.4 x 10^5) – (2.7 x 10^5) = 2.7 x 10^5 cells

Cells to be added in T75 flask = 2.70E+05 cells

Cell concentration = 7000 cells/ml

Cells to be added = 2.7 x 10^5 cells

Volume of cells to be taken = (2.7 x 10^5)/7000 = 38.5 ml

(This is higher than the volume of flask, you may need to take a more concentrated cell suspension)